The (partial) cost sheet for the single product manufactured at Vienna Company follows.

Direct labor(15 hours @ $25)$375

Variable overhead(15 hours @ $2) 30

Fixed overhead(15 hours @ $4) 60

The master budget level of production is 78,000 direct-labor hours, which is also the production volume used to compute the fixed overhead application rate. Other information available for operations over the past accounting period include the following.

Actual variable overhead incurred $ 128,000
Actual fixed overhead incurred 326,800
Direct labor efficiency variance 158,000 U
Variable overhead price variance 26,000

F

Required:
a. What was the variable overhead efficiency variance?
b. What was the fixed overhead price variance?
c. What was the fixed overhead production volume variance?

please show all of the steps, thank you very much. I have posted 2 times already, no one answered correctly.
Answer #1

a. Variable overhead price variance = (Actual hours* standard variable overhead rate per hour) - Actual variable overhead

$26000 = (Actual hours x 2) - 128000

Actual hours = 77000

Variable Overhead Efficiency Variance = Standard overhead rate x (Actual hours - Standard hours)

= (77000-78000)*2 = $2000 Favorable

b. Fixed overhead price variance = Actual fixed overhead incurred - budgeted fixed overhead

= $326800 - (78000*4)

= $14800 Unfavorable

c. Fixed overhead production volume variance = (Actual hours - budgeted hours)*Budgeted rate

= (77000-78000)*4

= $4000 Unfavorable